If displacement is dependent on time such that x=2t-2 then find out average speed up-to 4 seconds

Grade:12

3 Answers

Kanha singhania
6 years ago
Average velocity = total displacement / total time.Displacement varies as, x = 2t-2.Therefore total displacement= 2(4)-2 = 6mTotal time = 4secHence average velocity = 8/6 ~ 1.34
Shankarshan Dwivedi
3 years ago
Averageย  velocity =total displacement/total time taken.ย  Total displacement=2*4-2=6m. Total time taken =4sec. Average velocity =6/4=1.5.
Ayush_Deep
3 years ago
Dear friend the above two answers are wrong.
Average speed = total distance / total timeย 
Total time = 4 sec
Position of particle at t=0 is x= -2( initial)
Position of particle at t= 4 is x = 6(final)
Total distance traveled = final - initial
ย  ย  ย  ย  ย = 6-(-2)= 8
Hope understood
Good regards